OpenStudy (anonymous):

AP CALCULUS: The differentiable function g(x) has g(4) = 3 and g'(4) = -2. Let f(x) = x • g(x). The equation of the line tangent to f at x = 4 is... A) y = -8x + 35 B) y = 11x + 56 C) y = -2x + 20 D) y = -5x + 32 ***I know the answer is D, but I don't know how to get there. I realize you use product rule, but because the numbers are given it's confusing. Please explain?

OpenStudy (anonymous):

\[f(x)=xg(x)\] \[f'(x)=g(x)+xg'(x)\] \[f'(4)=g(4)+4g'(4)\] plug in the numbers, see what you get for the slope

OpenStudy (anonymous):

Thank you so much!